### v1

@@ Abstract
With the explosive growth of available training data, single-image 3D human modeling is ahead of a transition to a data-centric paradigm. A key to successfully exploiting data scale is to design flexible models that can be supervised from various heterogeneous data sources produced by different researchers or vendors. To this end, we propose a simple yet powerful paradigm for seamlessly unifying different human pose and shape-related tasks and datasets. Our formulation is centered on the ability – both at training and test time – to query any arbitrary point of the human volume, and obtain its estimated location in 3D. We achieve this by learning a continuous neural field of body point localizer functions, each of which is a differently parameterized 3D heatmap-based convolutional point localizer (detector). For generating parametric output, we propose an efficient post-processing step for fitting SMPL-family body models to nonparametric joint and vertex predictions. With this approach, we can naturally exploit differently annotated data sources including mesh, 2D/3D skeleton and dense pose, without having to convert between them, and thereby train large-scale 3D human mesh and skeleton estimation models that outperform the state-of-the-art on several public benchmarks including 3DPW, EMDB and SSP-3D by a considerable margin.
@@ 1 Introduction
Specifically, this localizer field’s input domain is the full 3D volume of the canonical human body and its output range is the parameter space of the modulated convolutional output layer (making the localizer field a hypernetwork [30]). By optimizing the neural field’s parameters, we obtain a smooth field of localizers (functions) over the human volume domain, improving robustness, and knowledge sharing. In the paper, we investigate various design choices related to this architecture, including positional encodings. This architecture results in consistent and accurate skeletal joint predictions along with smooth nonparametric mesh output, whose format is chosen by the user at test time, see Fig. 1. Since a parametric representation (for example based on SMPL [59]) is often preferred for further downstream processing, as the second contribution of this paper, we derive a simple and efficient algorithm for fitting SMPL-family body models to our non-parametric predictions. This involves alternating between estimating global body part orientations through the Kabsch algorithm [41], and solving for the shape coefficients in a regularized linear least squares formulation. This converges after as few as ca. 2–4 rounds, and is well-suited for GPU acceleration.
@@ 2 Related Work
Some works [49, 62] predict further keypoints on the surface beyond the skeleton, noting that this carries richer shape information. Ma et al. [62] predict a sparse set of virtual markers and approximate the rest through convex interpolation. Starting with Lin et al.’s pioneering work on the Mesh Transformer [57], nonparametric mesh estimation has been a successful line of recent research [17, 115, 20, 18, 48, 70, 61]. The main motivation for this is better preservation of alignment in contrast to parametric estimation, which often suffers from mean-shape bias [89]. In contrast to our work, these Transformer-based approaches feed a static set of keypoints into their architecture throughout training, which then exchange information through self-attention layers, and require a coarse to fine strategy and complex design for efficient processing. In contrast, our architecture is more lightweight, and can directly produce predictions for all possible points, without having to interpolate them from some other set of points.
@@ 3.1 Point Localizer Network
Given an RGB image crop of a person, we first extract feature maps through an off-the-shelf backbone at stride . We then follow prior work [85] in skeletal 3D pose estimation, and produce metric-scale truncation-robust volumetric heatmaps , as well as 2D pixel-space heatmaps through a 11 convolutional layer. The heatmaps are decoded to coordinates via soft-argmax [50, 74, 97], and the resulting 2D and root-relative 3D points are combined to an absolute camera-space estimate, including estimates for truncated body parts as in [85]. Additionally, we predict an uncertainty map , which will be averaged to an uncertainty scalar per point using weights from .
@@ 3.2 Neural Localizer Field
[…]
where the posed point is determined based on image features by . Although neural fields are typically used to predict points or vectors, here we use them to predict localizer functions . This way, we represent information on a continuous domain, while appropriately modeling both smooth structure and fine-grained details of a signal. Specifically, is implemented as a single-layer convolutional head whose weights are predicted as a function of the canonical point we want to localize. The function is realized as a feed-forward neural net, taking three canonical coordinates of as input and producing parameters () for ’s convolutional layer to control which point gets predicted. Here the stands for the backbone feature dimensionality plus a bias value, and are the total number of output channels consisting of depth channels of the volumetric heatmap , one 2D heatmap and one uncertainty map . In summary, given the image features and a canonical query point , the following computation takes place:
[…]
after which the 3D root-relative and the 2D predictions are fused into a camera-space output . For the technical details of this, we refer to [85]. The Neural Localizer Field’s architecture consists of a positional encoding part and an MLP with GELU activation (see the appendix for the details of the MLP layers).
@@
[…]
@@
Data in HPE contains a wide variety of annotation types including parametric meshes, 3D keypoints, 2D keypoints and DensePose samples. Thanks to our formulation described above, we can directly consume this data, since the zoo of annotations consists of subsets of points in the canonical volume. For examples that are annotated with a parametric mesh model, we sample 640 points uniformly over the surface and 384 points from the interior of the human volume. The sampled points are unique to each example of the batch. We supervise interior points densely to enable the network to predict new joint sets anywhere in the volume at test time. To deform interior points, we use scattered data interpolation [95, 8] to propagate SMPL deformations to the volume. For datasets annotated with 2D or 3D keypoints, the corresponding canonical positions need to be defined per joint. We use an approximate initialization per skeleton type, and finetune the precise canonical positions during the main training process. In case of DensePose annotations, we map the labels to the corresponding SMPL surface points in the canonical pose.
@@ 3.3 Efficient Body Model Fitting in Post-Processing
Above, we introduced our method to nonparametrically localize any human point from an image, in 3D. While nonparametric prediction has its advantages (e.g., closer fits to the image data), in some cases a parametric representation (joint rotations and shape vector ) is preferable due to its compactness, and disentanglement of pose and shape. We therefore develop a fast model-fitting algorithm to solve for pose and shape parameters that closely reproduce our predicted nonparametric vertices. Our algorithm alternates between two steps for a low number of iterations (3), followed by an adjustment step. We first independently fit global orientations per body part by applying a weighted Kabsch algorithm. Then, keeping the orientations fixed, we solve for and the translation vector via linear least squares, since SMPL’s vertex positions are linear in the shape parameters. Finally, we re-estimate the body-part rotations in one pass along the kinematic tree, this time non-independently but anchoring the rotations at the parent joint. This helps when the fitted yields slightly different bone lengths compared to the nonparametric estimate. For improved efficiency, it is possible to consider only a subset of the vertices when fitting. Further, given that our model estimates per-point uncertainties, these can be used in a weighted variant of the above algorithm. Splitting the shape estimation task into a nonparametric point localization and a fitting step has the further benefit that it is easy to share the body shape parameters for multiple observations of the same person (multi-view and temporal use cases) during the least-squares regression. More details and the pseudocode of the algorithm are given in the appendix.
@@ 4 Experiments
We extensively evaluate our method on a variety of benchmark test datasets: 3DPW [107] and EMDB [43] for SMPL body, AGORA [79] for SMPL-X, SSP-3D [89] for SMPL focusing on body shape, as well as Human3.6M [37], MPI-INF-3DHP [64] and MuPoTS-3D [66] for 3D skeletons.
Our goal is to enable large-scale multi-dataset training from heterogeneous annotation sources in order to train strong models. To show that our formulation is effective at this, we assemble a large meta-dataset for use in training. Due to space constraints we provide the detailed dataset description in the appendix. For parametric meshes we include the datasets [79, 7, 5, 32, 46, 102, 22, 33, 14, 122, 114, 10, 112, 116, 81, 9, 105], with points sampled directly from each dataset’s native body model (one of SMPL, SMPL-X or SMPL-H, neutral or gendered, as provided) without a need for any conversions. For 3D skeleton annotations, we use [37, 64, 66, 117, 123, 40, 52, 75, 108, 82, 31, 27, 124, 104, 76, 103, 21, 4, 26, 24, 25, 3, 44, 15]. Several of these have custom skeletons that are nontrivial to convert to the SMPL family, preventing existing 3D HPS methods from exploiting these rich data sources. For 2D keypoint annotations we use [34, 38, 106, 2, 109, 23] and further use DensePose datasets [72, 29] which provide pairs of SMPL surface points and corresponding pixel locations. Integrating all these datasets would be a great challenge if we had to choose one specific representation. Instead, we can easily train using the strategy described in Sec. 3.2, we simply map each annotation convention to the canonical human volume and learn a continuous localizer field.
@@
NLF outperforms all baselines on EMDB (Tab. 4), which has challenging outdoor sequences. The improvement is drastic, we obtain an MPJPE of 68.4 mm compared to the second-best BEDLAM-CLIFF (98.0 mm). Remarkably, we outperform the temporal method WHAM (79.7 mm) although we work frame-by-frame. 3DPW results confirm the same (Tab. 2). (For reference, the standard error of the mean under normal assumptions for 3DPW MPJPE is 0.06 mm.) We attribute this performance to the fact that we can natively train on many sources without error-prone conversions. (For better comparability to temporal methods, we also provide NLF results using a (non-learned) 5-frame temporal smoothing filter.) On AGORA (Tab. 5), we outperform all prior published works in SMPL-X prediction. Even though we do not specially target facial and hand keypoints, we still obtain the second best scores for these. With the same model, we also obtain excellent performance on the skeleton estimation benchmarks Human3.6M, MPI-INF-3DHP and MuPoTS-3D (Tab. 6). This evaluation is easy with NLF, as we can pick at test time which points of the human volume to predict.
@@
| SSP3D mIoU | 3DPW P-MVE | EMDB P-MVE | 3DPW MPJAE | |
| Plain XYZ coordinates | 0.777 | 54.0 | 57.7 | 16.9 |
| Global point signature (GPS) | 0.789 | 54.2 | 57.5 | 15.8 |
| Learnable Fourier features | 0.804 | 52.8 | 57.0 | 15.4 |
| Learnable Fourier (GPS init.) | 0.812 | 52.8 | 56.4 | 15.3 |
@@
| NLF-S (5-crop) | 107.8 | 68.4 | 105.7 | 68.3 | 99.2 | 62.9 | 27.6 | 40.7 | 40.6 | 97.2 | 62.8 | 29.3 | 36.8 | 37.0 |
| NLF-L (5-crop) | 99.2 | 63.3 | 97.2 | 63.0 | 91.3 | 58.2 | 25.8 | 36.6 | 37.0 | 89.4 | 58.0 | 27.3 | 33.1 | 33.7 |
@@ 5 Conclusion
We proposed Neural Localizer Fields to upgrade human pose estimation architectures from finite and fixed joint sets to predicting any point in a continuous canonical human volume. In a functional approach, we proposed to store point-localizer functions in a neural field, as opposed to storing explicit convolutional weights for every joint the network can predict. This enabled us to train a large-scale generalist pose and shape estimation model, without any effort on relabeling the data sources to one format. Instead, we could supervise our model with any points that happened to be annotated for particular training examples. Besides this training advantage, we can choose at test time, which point in the volume to predict, allowing to output any format we wish. To make our output even more useful in downstream applications, we proposed an efficient fitting algorithm to obtain SMPL parameters. We demonstrated that our model trained on several datasets with different formats achieves state-of-the-art performance with significant improvement over baselines on several benchmarks. Our model exhibits good performance on outdoor, indoor and synthetic data, regardless of the skeleton/mesh format used for different benchmarks. With the recent trend of more datasets from different sources being available, NLF paves the way to leverage these elegantly and efficiently.

### v2

@@ Abstract
With the explosive growth of available training data, single-image 3D human modeling is ahead of a transition to a data-centric paradigm. A key to successfully exploiting data scale is to design flexible models that can be supervised from various heterogeneous data sources produced by different researchers or vendors. To this end, we propose a simple yet powerful paradigm for seamlessly unifying different human pose and shape-related tasks and datasets. Our formulation is centered on the ability – both at training and test time – to query any arbitrary point of the human volume, and obtain its estimated location in 3D. We achieve this by learning a continuous neural field of body point localizer functions, each of which is a differently parameterized 3D heatmap-based convolutional point localizer (detector). For generating parametric output, we propose an efficient post-processing step for fitting SMPL-family body models to nonparametric joint and vertex predictions. With this approach, we can naturally exploit differently annotated data sources including mesh, 2D/3D skeleton and dense pose, without having to convert between them, and thereby train large-scale 3D human mesh and skeleton estimation models that considerably outperform the state-of-the-art on several public benchmarks including 3DPW, EMDB, EHF, SSP-3D and AGORA.
@@ 1 Introduction
Specifically, this localizer field’s input domain is the full 3D volume of the canonical human body and its output range is the parameter space of the modulated convolutional output layer (making the localizer field a hypernetwork [33, 46, 13]). By optimizing the neural field’s parameters, we obtain a smooth field of localizers (functions) over the human volume domain, improving robustness, and knowledge sharing. In the paper, we investigate various design choices related to this architecture, including positional encodings. The architecture results in consistent and accurate skeletal joint predictions along with smooth nonparametric mesh output, whose format is chosen by the user at test time, see Fig. 1. Since a parametric representation (for example based on SMPL [65]) is often preferred for further downstream processing, as the second contribution of this paper, we derive a simple and efficient algorithm for fitting SMPL-family body models to our non-parametric predictions. This involves alternating between estimating global body part orientations through the Kabsch algorithm [44], and solving for the shape coefficients in a regularized linear least squares formulation. This converges after as few as ca. 2–4 rounds, and is well-suited for GPU acceleration.
@@ 2 Related Work
Some works [53, 68] predict further keypoints on the surface beyond the skeleton, noting that this carries richer shape information. Ma et al. [68] predict a sparse set of virtual markers and approximate the rest through convex interpolation. Nonparametric mesh estimation has been a successful line of recent research [19, 122, 62, 63, 22, 20, 52, 76, 67], mainly due to better preservation of alignment in contrast to parametric estimation, which often suffers from mean-shape bias [96]. In contrast to our work, the prominent Transformer-based approaches of this category [62, 19, 63] feed a static set of keypoints into their architecture throughout training, which then exchange information through self-attention layers, and require a coarse to fine strategy and complex design for efficient processing. Our architecture is more lightweight, and can directly produce predictions for all possible points, without having to interpolate them from some other set of points.
@@ 3.1 Point Localizer Network
Given an RGB image crop of a person, we first extract feature maps through an off-the-shelf backbone at stride . We then follow the MeTRAbs [92, 93] architecture in skeletal 3D pose estimation, and produce metric-scale truncation-robust volumetric heatmaps , as well as 2D pixel-space heatmaps through a 11 convolutional layer. The heatmaps are decoded to coordinates via soft-argmax [54, 80, 104], and the resulting 2D and root-relative 3D points are combined to an absolute camera-space estimate, including estimates for truncated body parts. Additionally, we predict an uncertainty map , which will be averaged to an uncertainty scalar per point using weights from .
@@ 3.2 Neural Localizer Field
where the posed point is determined based on image features by . Although neural fields are typically used to predict points or vectors, here we use them to predict localizer functions . This way, we represent information on a continuous domain, while appropriately modeling both smooth structure and fine-grained details of a signal. Specifically, is implemented as a single-layer convolutional head whose weights are predicted as a function of the canonical point we want to localize. The function is realized as an MLP, taking three canonical coordinates of as input and producing parameters () for ’s convolutional layer to control which point gets predicted (cf. hypernetworks). Here the is the backbone feature dimensionality plus one bias, and are the total number of output channels which are depth channels of the volumetric heatmap , one 2D heatmap and one uncertainty map . In summary, given the image features and a canonical query point , the following computation takes place:
after which the 3D root-relative and the 2D predictions are fused into a camera-space output following [92]. The localizer field’s architecture consists of a positional encoding part and an MLP with GELU activation (see the appendix for the details of the MLP layers).
@@
In practice, we often do not have the full volume annotation available for all as in (8), and it would be too expensive to evaluate the integral for every image. Since we train on a mixture of datasets, each training example can contain a different annotated point set depending on the dataset the example comes from. Hence, we turn the integral (8) into a discrete sum over the available dataset-specific locations . For training examples with parametric body annotations, we perform Monte Carlo approximation of the integral by random sampling of points.
@@
Data in HPE contains a wide variety of annotation types including parametric meshes, 3D keypoints, 2D keypoints and DensePose samples. Thanks to our formulation described above, we can directly consume this data, since the zoo of annotations consists of subsets of points in the canonical volume. For examples that are annotated with a parametric mesh model, we sample 640 points uniformly over the surface and 384 points from the interior of the human volume. The sampled points are unique to each example of the batch. We supervise interior points densely to enable the network to predict new joint sets anywhere in the volume at test time. To deform interior points, we use scattered data interpolation [102, 9] to propagate SMPL deformations to the volume. For datasets annotated with 2D or 3D keypoints, the corresponding canonical positions need to be established per joint. We use an approximate initialization per skeleton type (see appendix), and finetune the precise canonical positions during the main training process. In case of DensePose annotations, we map the labels to the corresponding SMPL surface points in the canonical pose.
@@ 3.3 Efficient Body Model Fitting in Post-Processing
Above, we introduced our method to nonparametrically localize any human point from an image, in 3D. While nonparametric prediction has its advantages (e.g., closer fits to the image data), in some cases a parametric representation (joint rotations and shape vector ) is preferable due to its compactness, and disentanglement of pose and shape. We therefore develop a fast algorithm to solve for pose and shape parameters that closely reproduce our predicted nonparametric vertices.
Our algorithm alternates between two steps for a small number of iterations (3), followed by an adjustment step. We first independently fit global orientations per body part by applying a weighted Kabsch algorithm. Then, keeping the orientations fixed, we solve for and the translation vector via linear least squares, since SMPL’s vertex positions are linear in the shape parameters. Finally, we refine the part rotations in one pass along the kinematic tree, anchoring the rotations at the parent joint. For improved efficiency, it is possible to consider only a subset of the vertices when fitting. Further, given that our model estimates per-point uncertainties, these can be used in a weighted variant of the above algorithm. Splitting the shape estimation task into a nonparametric point localization and a fitting step has the further benefit that it is easy to share the body shape parameters for multiple observations of the same person (multi-view and temporal use cases) during the least-squares regression. More details and the pseudocode of the algorithm are given in the appendix.
@@ 4 Experiments
We extensively evaluate our method on a variety of benchmarks: 3DPW [114] and EMDB [47] for SMPL body, AGORA [85] and EHF [87] for SMPL-X, SSP-3D [96] for SMPL focusing on body shape, as well as Human3.6M [40], MPI-INF-3DHP [70] and MuPoTS-3D [72] for 3D skeletons.
Our goal is to enable large-scale multi-dataset training from heterogeneous annotation sources in order to train strong models. To show that our formulation is effective at this, we assemble a large meta-dataset for training, extending PosePile111https://github.com/isarandi/posepile, introduced in [93]. We provide dataset details in the appendix. For parametric meshes we include the datasets [85, 8, 6, 35, 50, 109, 24, 36, 16, 130, 121, 11, 119, 123, 88, 10, 112], with points sampled directly from each dataset’s native body model (one of SMPL, SMPL-X or SMPL-H, neutral or gendered, as provided) without a need for any conversions. For 3D skeleton annotations, we use [40, 70, 72, 124, 131, 43, 57, 81, 115, 89, 34, 29, 132, 111, 82, 110, 23, 5, 28, 26, 27, 4, 48, 17]. Several of these have custom skeletons that are nontrivial to convert to the SMPL family, preventing existing 3D HPS methods from exploiting these rich data sources. For 2D keypoint annotations we use [37, 41, 113, 2, 116, 25] and further use DensePose datasets [78, 31] which provide pairs of SMPL surface points and corresponding pixel locations. Integrating all these datasets would be a great challenge if we had to choose one specific representation. Instead, we can easily train using the strategy described in Sec. 3.2, we simply map each annotation convention to the canonical human volume and learn a continuous localizer field.
@@
| SSP3D mIoU | 3DPW P-MVE | EMDB P-MVE | 3DPW MPJAE | |
| Plain XYZ coordinates | 0.777 | 54.0 | 57.7 | 16.9 |
| Global point signature (GPS) | 0.789 | 54.2 | 57.5 | 15.8 |
| Learnable Fourier features | 0.804 | 52.8 | 57.0 | 15.4 |
| Learnable Fourier (GPS init.) | 0.812 | 52.8 | 56.4 | 15.3 |
@@
NLF outperforms all baselines on EMDB (Tab. 4), which has challenging outdoor sequences. The improvement is drastic, we get an MPJPE of 68.4 mm compared to the second-best (98.0 mm). Remarkably, we outperform the temporal method WHAM (79.7 mm) although we work frame-by-frame. 3DPW results confirm the same (Tab. 2). (For reference, the standard error of the mean under normal assumptions for 3DPW MPJPE is 0.06 mm.) We attribute this performance to the fact that we can natively train on many sources without error-prone conversions. (For better comparability to temporal methods, we also provide NLF results using a (non-learned) 5-frame temporal smoothing filter.) On AGORA (Tab. 5), we outperform all prior works in SMPL-X prediction. Even though we do not specially target facial and hand keypoints, we obtain the second best scores for these. We also achieve state-of-the-art body results on EHF (Tab. 7). The higher hand error is due to few detailed hands in the training data. We obtain excellent performance on the skeleton estimation benchmarks Human3.6M, MPI-INF-3DHP and MuPoTS-3D as well (Tab. 6). This evaluation is easy with NLF, as we can pick at test time which points of the body to predict.
@@
| Multi-HMR [3] | 102.0 | 63.4 | 101.8 | 64.1 | 95.9 | 59.6 | 27.7 | 40.2 | 40.9 | 95.7 | 60.3 | 29.2 | 38.1 | 39.0 |
| NLF-S | 108.6 | 67.9 | 106.3 | 67.4 | 102.1 | 63.8 | 29.5 | 42.7 | 42.8 | 99.9 | 63.4 | 31.7 | 38.7 | 39.2 |
| NLF-L | 98.6 | 62.1 | 96.6 | 61.9 | 92.7 | 58.4 | 27.0 | 37.9 | 38.1 | 90.8 | 58.2 | 28.5 | 34.4 | 34.9 |
| Method | Human3.6M | MPI-INF-3DHP | MuPoTS-3D | |||
| MPJPE | P-MPJPE | PCK | AUC | MPJPE | PCK-detected | |
| MeTRAbs-ACAE-S [93] | 40.2 | 31.1 | 96.3 | 58.7 | 57.9 | 94.7 |
| MeTRAbs-ACAE-L [93] | 36.5 | 27.8 | 97.1 | 60.1 | 55.4 | 95.4 |
| NLF-S | 40.4 | 30.6 | 96.6 | 57.9 | 59.9 | 94.7 |
| NLF-L | 39.7 | 28.5 | 97.5 | 61.0 | 54.9 | 95.5 |
| Method | MVE | P-MVE | P-MPJPE | ||||||
| Full-body | Hands | Face | Full-body | Body | Hands | Face | Body (14j) | Hands | |
| HybrIK-X [56] | 121.4 | 52.1 | 41.9 | 59.8 | 50.0 | 17.6 | 8.1 | 60.8 | 17.9 |
| OSX [61] | 70.8 | 53.7 | 26.4 | 48.7 | – | 15.9 | 6.0 | – | – |
| PyMAF-X [129] | 64.9 | 29.7 | 19.7 | 50.2 | 44.8 | 10.2 | 5.5 | 52.8 | 10.3 |
| SMPLer-X [12] | 62.4 | 47.1 | 17.0 | 37.1 | – | 14.1 | 5.0 | – | – |
| Multi-HMR [3] | 42.0 | 28.9 | 18.0 | 28.2 | – | 10.8 | 5.3 | – | – |
| NLF-S | 39.7 | 49.7 | 15.2 | 30.9 | 30.4 | 24.5 | 7.4 | 32.8 | 28.2 |
| NLF-S +fit | 40.0 | 49.7 | 15.4 | 30.4 | 29.7 | 24.0 | 6.5 | 32.0 | 25.4 |
| NLF-L | 36.3 | 43.2 | 13.5 | 25.8 | 24.8 | 21.3 | 6.7 | 26.3 | 25.3 |
| NLF-L +fit | 36.4 | 43.0 | 13.9 | 26.0 | 24.4 | 20.7 | 6.3 | 26.1 | 21.8 |
| SSP-3D | H36M | 3DHP | 3DPW | 3DPW | EMDB | EMDB | |
| Method / data | PVE-T-SC | MPJPE | PCK | P-MPJPE | P-MVE | P-MPJPE | P-MVE |
| NLF-S, real datasets only | 11.9 | 43.0 | 96.0 | 40.4 | 55.5 | 47.8 | 59.3 |
| NLF-S, synthetic datasets only | 10.0 | 48.0 | 95.4 | 39.2 | 53.5 | 45.7 | 57.9 |
| NLF-S, all datasets | 9.9 | 41.6 | 96.3 | 38.4 | 53.1 | 45.8 | 57.2 |
Effect of datasets
Assessing the contribution of each individual dataset is computationally infeasible, nonetheless, to validate that NLF benefits from diverse supervision, we partition the data into real and synthetic subsets. As Tab. 8 shows, best results are achieved when all datasets are used in training.
Real-time inference
NLF-S has a batched throughput of 410 fps and unbatched throughput of 79 fps on an Nvidia RTX 3090 GPU. For NLF-L these are 109 fps and 41 fps respectively.
@@ 5 Conclusion
We proposed Neural Localizer Fields to upgrade 3D human pose estimators from fixed joint sets to predicting any point of the body surface and volume. We store point-localizer functions in a neural field, as opposed to storing explicit convolutional weights for every joint the network can predict. This enables us to train a large-scale generalist pose and shape estimation model, without any effort on relabeling all data to one format. Instead, we can supervise our model with any points that happen to be annotated for particular training examples. Besides this training advantage, we can choose at runtime which points in the volume to predict, allowing us to output any desired format. To make our output even more useful in downstream applications, we proposed an efficient points-to-parameters fitting algorithm to obtain SMPL parameters. Our model trained on several datasets with different formats achieves state-of-the-art performance with significant improvement over baselines on several benchmarks. Our model exhibits good performance on outdoor, indoor and synthetic data, regardless of the skeleton/mesh format used for different benchmarks. With the recent trend of more datasets from different sources being available, NLF paves the way to leverage these elegantly and efficiently.